Grade 11Physical Chemistry

If the velocity of the first orbit of hydrogen is x then the velocity of third orbit will be???

harshal pawar
8 Years agoGrade 11

1 Answer

Dwij
8 Years ago
V = 2.18 x 10^8 Z/n cm/ sec = x V for 3rd shell = (2.18 x10^8 )/3 = x/3V = velocity of electron, Z = at. No. (in this case 1) and n = no.of shell